OpenStudy (dmezzullo):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5. y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (dmezzullo):

@terryluce06

OpenStudy (dmezzullo):

@sue101

OpenStudy (ajprincess):

The equation of a line with the slope m and passing through a point (x_1,y_1) is given by y-y_1=m(x-x_1) Does that help?

OpenStudy (dmezzullo):

Its easier for me if its not complicated

OpenStudy (dmezzullo):

@sue101 Can u help?

OpenStudy (anonymous):

Slope=(Y-y1)/(X-x1). 5=(Y+1)/(x+2). Y+1=5(x+2)

OpenStudy (dmezzullo):

huh?

OpenStudy (dmezzullo):

@terryluce06